the average age of 5 boys is 10, 2 new boys joined the group making the group average age 12, if the new boys are of the same age what is their actual age
Answer by TimothyLamb(4379)

x = age of new boy A
y = age of new boy B
---
(10*5 + x + y)/7 = 12
x = y
---
put the system of linear equations into standard form
---
x + y = 12*7 - 50
x - y = 0
---
x + y = 34
x - y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
solution:
x = age of new boy A = 17
y = age of new boy B = 17
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---